A 240-turn solenoid is 80.0cm long and carries a current of 0.30A. Calculate the magnitude of the magnetic field at the center of solenoid.


1
Expert's answer
2022-05-26T14:07:15-0400

Magnetic field inside solenoid is given by

"B=\mu_0 \frac{N}{L}I"

"B=4\pi*10^{-7}* \frac{240}{0.80}*0.30=1.13*10^{-4}\rm T"